# On complexity of restricted fragments of Decision DNNF

### v1

@@ Abstract
In the remaining part of the paper, we introduce a relaxed version of Structured Decision dnnf that we name Structured $\wedge_{d}$-fbdd. In particular, we explain why Decision dnnf is equivalent to $\wedge_{d}$-fbdd but their structured versions are distinct.) We demonstrate that this model is quite powerful for cnfs of bounded incidence treewidth: it has an fpt representation for cnfs that can be turned into ones of bounded primal treewidth by removal of a constant number of clauses (while for both $\wedge_{d}$-obdd and Structured Decision dnnf an xp lower bound is triggered by just two long clauses).
@@ 1.1 The main objective
Decomposable negation normal forms (dnnfs) [5] is a model for representation of Boolean functions that is of a landmark importance in the area of (propositional) Knowledge compilation (kc). The difference of the model from deMorgan circuits is that all the $\wedge$ gates are decomposable meaning, informally, that two distinct inputs of the gate cannot be reached by the same variable. We will refer to such gates as $\wedge_{d}$. Another way to define dnnfs is as a Non-deterministic read-once branching program ($1$-nbp) equipped with $\wedge_{d}$ gates. This latter view is useful for understanding that (from the univariate perspective, that is when the size bounds depend on the number $n$ of input variables) dnnfs are not very strong compared to $1$-nbps since the latter can quasipolynomially simulate the former [14]. In particular, a class of Boolean functions requiring exponentially large $1$-nbps for its presentation requires an exponential lower bound for representation by dnnfs.
The strength of $\wedge_{d}$ becomes distinctive when we consider dnnfs from the multivariate perspective. In particular, cnfs of primal treewidth at most $k$ have fpt presentation as a dnnf [5] but, in general requires an xp-sized representation as a $1$-nbp. [15]. The fpt upper bound as above turned out to be robust in the sense that it is preserved by all known restrictions of dnnfs that retain $\wedge_{d}$-gates. In light of this, it is natural to investigate if fpt upper bounds hold for the incidence treewidth.
Among all known dnnf restrictions using $\wedge_{d}$ gates, there are two that are minimal in the sense that every other restriction is a generalization of one of them. One such minimal restriction is a Sentential Decision Diagram [6] that admits fpt presentation for a much more general class that cnfs of bounded incidence treewidth [2]. The other restriction is Decision dnnf [8, 11] where the parameterized complexity of representation of cnfs of bounded incidence treewidth is an open question formally stated below.
@@
To understand the reason of difficulty of Open Question 1, it is convenient to view Decision dnnfs as Free Binary Decision Diagrams (fbdds) equipped with $\wedge_{d}$ gates [1]. Throughout this paper, we will refer to this equivalent representation as $\wedge_{d}$-fbdd. This representation restricts the role of $\vee$ gates only to the representation of Shannon expansion. That is each $\vee$ gate is associated with a variable and represents a function $(x\wedge f)\vee(\neg x\wedge g)$ where $f$ and $g$ are some Boolean functions not depending on $x$. This restriction enables an easy guess of the variables of variables but it is not clear how to handle the guessing of satisfaction/non-satisfaction of clauses.
@@
The main objective of the research presented in this paper is to make a progress towards resolution of Open Question 1. Before presenting our results, we state reasons why the question is important to resolve.
@@ 1.2 Motivation
It has been observed in [8] that the trace of several well known dpll based model counters (##\##-sat solvers) using component analysis can be represented as a Decision dnnf. It is well known that the number of satisfying assignments of a cnf of bounded incidence treewidth can be efficiently computed [17]. On the other hand, it is not clear whether the efficient computation can be carried out by the dppl solvers. Resolving Open Question 1 will provide an important insight in this direction.
Another aspect of the motivation is the relationship between $\wedge_{d}$-fbdd and proof complexity. In particular, there is a close similarity between $\wedge_{d}$-fbdd and regular resolution as a well known fbdd-based representation of regular resolution implicitly uses decomposable conjunction. Indeed, to prove unsatisfiability of two variable disjoint cnfs, it is enough to prove unsatisfiability of just one of them and to completely ignore the other. Importantly, it is an open question whether unsatisfiable cnfs of bounded incidence treewidth have fpt-sized regular resolution refutation. In the context of the similarity as specified above, it is reasonable to expect that understanding the power of $\wedge_{d}$-fbdd as related to cnfs of bounded incidence treewidth would provide an insight for such a power of regular resolution. This connection is especially intriguing as the question for regular resolution is a restriction of the same question for the full-fledged resolution, a well known open problem in the area of proof complexity [4]
1.3 Our results and structure of the paper
In this paper we tackle the above question by restricting it to special cases of $\wedge_{d}$-fbdd. We consider two restrictions: obliviosness and structuredness.
Arguably, the best known restricted class of fbdd is Ordered Binary Decision Diagram (obdd). Applying this restriction to $\wedge_{d}$-fbdds, we obtain $\wedge_{d}$-obdds or, to put it differently, obdds equipped with $\wedge_{d}$ gates. Arguing as in [8], we conclude that $\wedge_{d}$-obdd represents the trace of a dblp model counter with component analysis and fixed variable ordering
While $\wedge_{d}$-obdd are known to have fpt representation [12] (as noted in [3] the constructiom in [12] is, in fact, structured $\wedge_{d}$-obdd), it is shown in [3] that representation of cnfs of bounded incidence treewidth requires an xp-sized representation as $\wedge_{d}$-obdds.
In this paper, we restate the proof of [3] by explicitly stating a generic approach to proving lower bounds of $\wedge_{d}$-obdd as a generalization of such an approach for obdd. Using this generic approach we demonstrate an exponential separation between fbdd and $\wedge_{d}$-obdd. In particular, we introduce a class of cnfs representable by poly-sized fbdd and prove an exponential lower bound for representation of this class of cnfs as $\wedge_{d}$-obdds.
Our last result for $\wedge_{d}$-obdds is an exponential separation between $\wedge_{d}$=obdds and obdds. In the context of sat solvers this separation means that the use of component analysis in presence of a fixed variable ordering can lead to an exponential runtime improvement. This situation contrasts with the dynamic variable ordering because of a well known quasipolynomial simulation of $\wedge_{d}$-textscfbdd by fbdd [1].
As stated above, the second restriction on $\wedge_{d}$-fbdds considered in this paper is structuredness. In the context of cnfs of bounded incidence treewidth structured Decision dnff is a rather restrictive model. Indeed, it is shown in [3] that taking a monotone $2$-cnf of bounded primal treewidth and adding one clause with negative literals of all the variables requires structured Decision dnnfs representing such cnfs to be xp-sized in terms of their incidence treewidth. By weakening the structurality restriction, we introduce a more powerful model that we name structured $\wedge_{d}$-fbdds.
The starting point of the construction is that transformation from Decision dnnf into $\wedge_{d}$-fbbd leads to an encapsulation effect: a $\vee$ node and two $\wedge_{d}$ nodes implementing Shannon decomposition are replaced by a single decision node. The structured $\wedge_{d}$-fbdd requires that only the $\wedge_{d}$ nodes explicitly present in the model respect a vtree or to put it differently, no constraint is impossed on the $\wedge_{d}$ nodes ’hidden’ inside the decision nodes.
In the rest of the paper, Section 2 introduces the necessary technical background, Sections 3 and 4 present the results related to $\wedge_{d}$-obdd and the structured restrictions of Decision dnnf. Proofs of selected statements are postponed to the Appendix.
@@ 2.1 Assignments, rectangles, Boolean functions
We are now going to introduce an operation that is, conceptually, very similar to the Cartesian product. We thus slightly abuse the notation by referring to this operation as the Cartesian product. So, let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be sets of assignments with $\mathsf{var}(\mathcal{H}_{1})\cap\mathsf{var}\mathcal{H}_{2})=\emptyset$. We denote $\{{\bf a}\cup{\bf b}|{\bf a}\in\mathcal{H}_{1},{\bf b}\in\mathcal{H}_{2}\}$ by $\mathcal{H}_{1}\times\mathcal{H}_{2}$. For example, suppose that $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\},\{(x_{2},1)\}\}$ and $\mathcal{H}_{2}=\{\{(x_{3},1)\},\{(x_{3},0),(x_{4},0)\}\}$. Then $\mathcal{H}_{1}\times\mathcal{H}_{2}=\{\{(x_{1},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},0),(x_{3},0),(x_{4},0)\},\{(x_{2},1),(x_{3},1)\},\{(x_{2},1),(x_{3},0),(x_{4},0)\}\}$. We note that for any set $\mathcal{H}$ of assignments, $\mathcal{H}\times\{\emptyset\}=\mathcal{H}$.
@@
Let $\mathcal{H}$ be a rectangle and let $Y\subseteq\mathsf{var}(\mathcal{H})$. We say that $\mathcal{H}$ breaks $Y$ if there is a pair $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of witnesses of $\mathcal{H}$ such that both $Y\cap\mathsf{var}(\mathcal{H}_{1})$ and $Y\cap var(\mathcal{H}_{2})$ are nonempty.
@@
We can naturally extend $\mathcal{H}_{1}\times\mathcal{H}_{2}$ to the case of more than two sets of assignments, that is $\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m}$ as e.g. $(\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m-1})\times\mathcal{H}_{m}$. The operation is clearly commutative that is we can permute the elements in any order. We also use the notation $\prod_{i\in[m]}\mathcal{H}_{i}$ (recall that $[m]=\{1,\dots,m\}$) We notice that if $I_{1},I_{2}$ is a partition of $[m]$ then $\prod_{i\in[m]}\mathcal{H}_{i}=\prod_{i\in I_{1}}\mathcal{H}_{i}\times\prod_{i\in I_{2}}\mathcal{H}_{i}$. Therefore, we observe the following.
@@
[…]
For example, if =$\mathcal{H}=\{\{(x_{1},0),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},1),(x_{3},0\}\}$ and ${\bf a}=\{(x_{1},1)\}$ then $\mathcal{H}|_{\bf a}=\{\{(x_{2},0),(x_{3},1)\},\{(x_{2},1),(x_{3},0)\}\}$.
@@
A literal is a variable $x$ or its negation $\neg x$. For a literal $\ell$, we denote its variable by $\mathsf{var}(\ell)$. Naturally for a set $L$ of literals $\mathsf{var}(L)=\{var(\ell)|\ell\in L\}$. $L$ is well formed if for every two distinct literals $\ell_{1},\ell_{2}\in L$, $\mathsf{var}(\ell_{1})\neq\mathsf{var}(\ell_{2})$.

### v2

@@ Abstract
We also study the complexity of Apply operation for $\wedge_{d}$-obdd. We demonstrate that, in general, the Apply operation leads to exponential blow up of the resulting model. We identify a special restricted case where the Apply operation can be carried out efficiently. In order to define the special case, we consider a subclass of $\wedge_{d}$-obdd embeddable into a linear order and introduce a novel concept of irregularity index measuring how far the model from an ordinary obdd. We demonstrate that the efficiency of Apply for the cases where two models are embeddable into the same order with low irregularity indices.
In the remaining part of the paper, we introduce a relaxed version of Structured Decision dnnf that we name Structured $\wedge_{d}$-fbdd. In particular, we explain why Decision dnnf is equivalent to $\wedge_{d}$-fbdd but their structured versions are distinct. We demonstrate that this model is quite powerful for cnfs of bounded incidence treewidth: it has an fpt representation for cnfs that can be turned into ones of bounded primal treewidth by removal of a constant number of clauses (while for both $\wedge_{d}$-obdd and Structured Decision dnnf an xp lower bound is triggered by just two long clauses).
@@ 1.1 The main objective
Decomposable negation normal forms (dnnfs) [7] is a model for representation of Boolean functions that is of a landmark importance in the area of (propositional) Knowledge compilation (kc). The difference of the model from deMorgan circuits is that all the $\wedge$-gates are decomposable meaning, informally, that two distinct inputs of the gate cannot be reached by the same variable. We will refer to such gates as $\wedge_{d}$-gates. Another way to define dnnfs is as a Non-deterministic read-once branching program ($1$-nbp) equipped with $\wedge_{d}$-nodes (in the context of branching programs, we refer to gates as nodes). This latter view is useful for understanding that (from the univariate perspective, that is when the size bounds depend on the number $n$ of input variables) dnnfs are not very strong compared to $1$-nbps since the latter can quasipolynomially simulate the former [18]. In particular, a class of Boolean functions requiring for its presentation $1$-nbps exponentially large in the number of variables, also requires exponentially large dnnfs.
The strength of $\wedge_{d}$-gates becomes distinctive when we consider dnnfs from the multivariate perspective. In particular, cnfs of primal treewidth at most $k$ have fpt presentation as a dnnf [7] but, in general require an xp-sized representation as a $1$-nbp. [19]. The fpt upper bound as above turned out to be robust in the sense that it is preserved by all known restrictions of dnnfs that retain $\wedge_{d}$-gates. In light of this, it is natural to investigate if fpt upper bounds hold for the incidence treewidth.
Among all known dnnf restrictions using $\wedge_{d}$-gates, there are two that are minimal in the sense that every other restriction is a generalization of one of them. One such minimal restriction is a Sentential Decision Diagram [8] that admits fpt presentation for a much more general class than cnfs of bounded incidence treewidth [4]. The other restriction is Decision dnnf [10, 14] where the parameterized complexity of representation of cnfs of bounded incidence treewidth is an open question formally stated below.
@@
To understand the reason of difficulty of Open Question 1, it is convenient to view Decision dnnfs as Free Binary Decision Diagrams (fbdds) equipped with $\wedge_{d}$-nodes [3]. Throughout this paper, we will refer to this equivalent representation as $\wedge_{d}$-fbdd. This representation clearly illustrates that the $\vee$-gates of Decision dnnf are only capable to ‘guess’ the value of variables associated with them. However, it is not clear how to handle the guessing of satisfaction/non-satisfaction of clauses.
@@
The main objective of the research presented in this paper is to make a progress towards resolution of Open Question 1. In the next subsection of the Introduction, we state our results, the motivation is discussed in Subsection 1.3, the structure of the rest of the paper is discussed in Subsection 1.4.
1.2 Our results
In this paper we tackle the Open Question 1 by restricting it to special cases of $\wedge_{d}$-fbdd. We consider two restrictions: obliviosness and structuredness.
Arguably, the best known restricted class of fbdd is Ordered Binary Decision Diagram (obdd). Applying this restriction to $\wedge_{d}$-fbdds, we obtain $\wedge_{d}$-obdds or, to put it differently, obdds equipped with $\wedge_{d}$-nodes. Arguing as in [10], we conclude that $\wedge_{d}$-obdd represents the trace of a dpll model counter with component analysis and fixed variable ordering.
While $\wedge_{d}$-obdd are known to have fpt representation [15] parameterized by the primal treewidth of the input cnf, (as noted in [5] the construction in [15] is, in fact, structured $\wedge_{d}$-obdd), it is shown in [5] that representation of cnfs parameterized by the bounded incidence treewidth requires an xp-sized representation as $\wedge_{d}$-obdds.
In this paper we introduce a generic methodology for proving lower bounds for $\wedge_{d}$-obdds. The key part of this methodology is a novel concept of an ($\wedge_{d}$-obdd-) fooling set $\mathcal{F}$ of assignments defined for a Boolean function $f$ and a permutation $\pi$ of $\mathsf{var}(f)$, the set of variables of $f$. We demonstrate that if $B$ is a $\wedge_{d}$-obdd representing function $f$ and respecting the order $\pi$ then $|B|\geq|\mathcal{F}|$. Thus proving large lower bounds reduces to establishing existence of a large fooling set.
We use the above methodology to obtain two lower bounds for $\wedge_{d}$-obdds. First, we restate the xp lower bound proof of [5]. Second, we demonstrate an exponential separation between fbdd and $\wedge_{d}$-obdd. In particular, we introduce a class of cnfs representable by poly-sized fbdd and prove an exponential lower bound for representation of this class of cnfs as $\wedge_{d}$-obdds.
We also demonstrate an exponential separation between $\wedge_{d}$-obdds and obdds.
An important property of obdd is an efficient implementation of Apply (conjunction) operation. In particular, let $B_{1}$ and $B_{2}$ be two obdds over the same set of variables and respecting the same order that represent functions $f(B_{1})$ and $f(B_{2})$ respectively. Then the function $f(B_{1})\wedge f(B_{2})$ can be represented by an obdd of size $|B_{1}|\times|B_{2}|$. It is thus natural to study whether $\wedge_{d}$-obdd allows efficient Apply operation. We demonstrate that in general this is not the case. In particular, we show that vertical and horizontal lines of an $n\times n$ grid can be represented as linear sized $\wedge_{d}$-obdd respecting the same ‘dictionary’ order. On the other hand, their conjunction is a ‘fully-fledged’ grid and it requires an exponential-sied representation even as a dnnf, let alone an $\wedge_{d}$-obdd.
In light of the above, it is natural to study a restriction of $\wedge_{d}$-obdd that enables efficient implimentation of Apply We introduce a notion of embeddability of a model into a linear order that can be seen as a weak form of structuredness. For Boolean circuits, this notion is known under the name of compatible orders [1], for probabilistic circuits, a similar idea of contiguity was used in [26] for a similar purpose of circuits multiplication.
We introduce a novel notion of an irregularity index that, intutively speaking, shows how far an $\wedge_{d}$-obdd respecting an order $\pi$ is from an obdd embeddable in the same order. We demonstrate that if $B_{1}$ and $B_{2}$ are $\wedge_{d}$-obdds respecting the same order with the respective irregularity indices $k_{1}$ and $k_{2}$ then $f(B_{1})\wedge f(B_{2})$ can be represented as a $\wedge_{d}$-obdd of size $(|B_{1}|\times|B_{2}|)^{O(k_{1}+k_{2})}$. We believe that the notion of irregularity idex gives rise to a new research direction of general interest in the area of kc. We provide the related discussion in the Conclusion section posing a number of conjectures and open questions.
As we stated above, the second restriction on $\wedge_{d}$-fbdds considered in this paper is structuredness. In the context of cnfs of bounded incidence treewidth, structured Decision dnnf is a rather restrictive model. Indeed, it is shown in [5] that taking a monotone $2$-cnf of bounded primal treewidth and adding one clause with negative literals of all the variables requires structured Decision dnnfs representing such cnfs to be xp-sized in terms of their incidence treewidth. By weakening the structurality restriction, we introduce a more powerful model that we name structured $\wedge_{d}$-fbdds.
The starting point of the construction is that transformation from Decision dnnf into $\wedge_{d}$-fbbd leads to an encapsulation effect: a $\vee$-gate and two $\wedge_{d}$-gates implementing Shannon decomposition are replaced by a single decision node. The structured $\wedge_{d}$-fbdd requires that only the $\wedge_{d}$-nodes explicitly present in the model respect a vtree or to put it differently, no constraint is imposed on the $\wedge_{d}$-gates ‘hidden’ inside the decision nodes.
@@ 1.3 Motivation
The motivation of the proposed results is double-fold. First, we make a progress towards resolution of Open Question 1. In this context, we demonstrate that $\wedge_{d}$-fbdd respecting a specific order cannot efficiently represent cnfs of bounded incidence treewidth. To put it differently, if $\wedge_{d}$-fbdd admits an fpt representation of cnfs of bounded incidence treewidth then this representation must heavily depend on absence of any static order of querying of variables.
We also introduce a novel concept of structured $\wedge_{d}$-fbdd, demonstrate its power as compared to the known models and propose understanding the complexity of its representation of cnfs of bounded incidence treewidth as the next step towards resolution of Open Question 1.
The second aspect of motivation is that our results for $\wedge_{d}$-obdd are interesting in their own right. Indeed, in the context of tracing of model counters, the model can be seen as tracing model counters with component analysis and a static ordering of variables. Our results lead to two conclusions about such solvers. First, the xp lower bound for cnfs of bounded incidence treewidth implies that the static ordering does not allow efficient handling of CNFs of bounded incidence treewidth.
Second, it is interesting to see the exponential separation between $\wedge_{d}$-obdd in contrast with quasi-polynomial simulation of $\wedge_{d}$-fbdd by fbdd [3]. This discrepancy demonstrates that, in presence of static ordering, the component analysis is way more important for efficiency of the solver than in case of a dynamical ordering. From the practical perspective, this conclusion means that, in presence of a static ordering, the time overhead spent to component analysis is highly worthwhile.
Furthermore, our results related to Apply operation for $\wedge_{d}$-obdd may be of interest for the area of kc in general beyond the formalisms considered in this paper. For example, the novel notion of irregularity index provides an efficient way of turning a restricted dnnf into a model of the corresponding class of read-once branching programs. Similar transformations are known for Decision dnffs and unrestricted dnnfs and result in a quasi-polynomial blow-up in the size of the resulting read-once branching programs [2, 12]. Moreover, as we show in this paper, the transformation from $\wedge_{d}$-obdd to obdd may even result in an exponential blow-up. However when the transformation is controlled by the irregularity index parameter, it becomes efficient if the value of the parameter is small.
1.4 Structure of the paper
The rest of the paper is organized as follows. The necessary terminology and the related statements are provided in Section 2. The xp lower bound and the separations for $\wedge_{d}$-obdd are provided in Section 3. The complexity of Apply operation for $\wedge_{d}$-obdd is discussed in Section 4. The structuredness applied to $\wedge_{d}$-fbdd is discussed in Section 5. The concluding discussion and the list of related open questions are provided in Section 6.
@@ 2.1 Assignments, rectangles, Boolean functions
We are now going to introduce an operation that is, conceptually, very similar to the Cartesian product. We thus slightly abuse the notation by referring to this operation as the Cartesian product. So, let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be sets of assignments with $\mathsf{var}(\mathcal{H}_{1})\cap\mathsf{var}(\mathcal{H}_{2})=\emptyset$. We denote $\{{\bf a}\cup{\bf b}|{\bf a}\in\mathcal{H}_{1},{\bf b}\in\mathcal{H}_{2}\}$ by $\mathcal{H}_{1}\times\mathcal{H}_{2}$. For example, suppose that $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\},\{(x_{2},1)\}\}$ and $\mathcal{H}_{2}=\{\{(x_{3},1)\},\{(x_{3},0),(x_{4},0)\}\}$. Then $\mathcal{H}_{1}\times\mathcal{H}_{2}=\{\{(x_{1},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},0),(x_{3},0),(x_{4},0)\},\{(x_{2},1),(x_{3},1)\},\{(x_{2},1),(x_{3},0),(x_{4},0)\}\}$. We note that for any set $\mathcal{H}$ of assignments, $\mathcal{H}\times\{\emptyset\}=\mathcal{H}$.
@@
Let $\mathcal{H}$ be a rectangle and let $Y\subseteq\mathsf{var}(\mathcal{H})$. We say that $\mathcal{H}$ breaks $Y$ if there is a pair $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of witnesses of $\mathcal{H}$ such that both $Y\cap\mathsf{var}(\mathcal{H}_{1})$ and $Y\cap\mathsf{var}(\mathcal{H}_{2})$ are nonempty.
@@
We can naturally extend $\mathcal{H}_{1}\times\mathcal{H}_{2}$ to the case of more than two sets of assignments, that is $\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m}$ as e.g. $(\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m-1})\times\mathcal{H}_{m}$. The operation is clearly commutative that is we can permute the elements in any order. We also use the notation $\prod_{i\in[m]}\mathcal{H}_{i}$ (recall that $m]=\{1,\dots,m\}$) We notice that if $I_{1},I_{2}$ is a partition of $[m]$ then $\prod_{i\in[m]}\mathcal{H}_{i}=\prod_{i\in I_{1}}\mathcal{H}_{i}\times\prod_{i\in I_{2}}\mathcal{H}_{i}$. Therefore, we observe the following.
@@
    Let $\bf g$} be an assignment and let $U$ be a set of variables. The projection $Proj({\bf g},U)$ is the assignment $\bf h$} such that $\mathsf{var}({\bf h})=\mathsf{var}({\bf g})\cap U$ and for each $v\in\mathsf{var}({\bf h})$, ${\bf h}(v)={\bf g}(v)$. Let $\mathcal{H}$ be a set of assignments and let $\bf a$} be an assignment such that $\mathsf{var}({\bf a})\subseteq\mathsf{var}(\mathcal{H})$. Then $\mathcal{H}|_{\bf a}=\{{\bf b}\setminus{\bf a}|{\bf b}\in\mathcal{H},{\bf a}\subseteq{\bf b}\}$. We call $\mathcal{H}|_{\bf a}$ the restriction of $\mathcal{H}$ by $\bf a$}. To extend the notion of restriction to the case where $\mathsf{var}({\bf a})$ is not necessarily a subset of $\mathsf{var}(\mathcal{H})$, we set $\mathcal{H}|_{\bf a}=\mathcal{H}_{Proj({\bf a},\mathsf{var}(\mathcal{H}))}$.
For example, if $\mathcal{H}=\{\{(x_{1},0),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},1),(x_{3},0\}\}$ and ${\bf a}=\{(x_{1},1)\}$ then $\mathcal{H}|_{\bf a}=\{\{(x_{2},0),(x_{3},1)\},\{(x_{2},1),(x_{3},0)\}\}$.
@@
A literal is a variable $x$ or its negation $\neg x$. For a literal $\ell$, we denote its variable by $\mathsf{var}(\ell)$. Naturally for a set $L$ of literals $\mathsf{var}(L)=\{\mathsf{var}(\ell)|\ell\in L\}$. $L$ is well formed if for every two distinct literals $\ell_{1},\ell_{2}\in L$, $\mathsf{var}(\ell_{1})\neq\mathsf{var}(\ell_{2})$.